Please help!
2. Convert the following equations into exponential form:
a. X = log38
b. -9 = log5Y
c. X= log6Y
d. 100 = log2Z

Answer by Alan3354(69443)

Please help!
2. Convert the following equations into exponential form:
a. X = log38
10^x = 38
------------
b. -9 = log5Y
10^(-9) = 5y
-----------
c. X= log6Y
10^x = 6y
---------
d. 100 = log2Z
10^100 = 2z